Methylamine, CH3NH2, has a Kb = 4.40 x 10-4 What is the pH of a 0.390 M solution of methylamine?
Answer #1

Lets write the dissociation equation of CH3NH2

CH3NH2 +H2O -----> CH3NH3+ + OH-

0.36 0 0

0.36-x x x

Kb = [CH3NH3+][OH-]/[CH3NH2]

Kb = x*x/(c-x)

4.4*10^-4 = x^2/(0.39-x)

since x can't be negative, the possible value of x is

x = 1.288*10^-2

so.[OH-] = x = 1.288*10^-2 M

we have below equation to be used:

pOH = -log [OH-]

= -log (1.288*10^-2)

= 1.89

we have below equation to be used:

PH = 14 - pOH

= 14 - 1.89

= 12.11

Answer: 12.11

Feel free to comment below if you have any doubts or if this answer do not work